%merlin.mbs apsrev4-1.bst 2010-07-25 4.21a (PWD, AO, DPC) hacked
%Control: key (0)
%Control: author (0) dotless jnrlst
%Control: editor formatted (1) identically to author
%Control: production of article title (0) allowed
%Control: page (1) range
%Control: year (0) verbatim
%Control: production of eprint (0) enabled
\begin{thebibliography}{44}%
\makeatletter
\providecommand \@ifxundefined [1]{%
 \@ifx{#1\undefined}
}%
\providecommand \@ifnum [1]{%
 \ifnum #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \@ifx [1]{%
 \ifx #1\expandafter \@firstoftwo
 \else \expandafter \@secondoftwo
 \fi
}%
\providecommand \natexlab [1]{#1}%
\providecommand \enquote  [1]{``#1''}%
\providecommand \bibnamefont  [1]{#1}%
\providecommand \bibfnamefont [1]{#1}%
\providecommand \citenamefont [1]{#1}%
\providecommand \href@noop [0]{\@secondoftwo}%
\providecommand \href [0]{\begingroup \@sanitize@url \@href}%
\providecommand \@href[1]{\@@startlink{#1}\@@href}%
\providecommand \@@href[1]{\endgroup#1\@@endlink}%
\providecommand \@sanitize@url [0]{\catcode `\\12\catcode `\$12\catcode
  `\&12\catcode `\#12\catcode `\^12\catcode `\_12\catcode `\%12\relax}%
\providecommand \@@startlink[1]{}%
\providecommand \@@endlink[0]{}%
\providecommand \url  [0]{\begingroup\@sanitize@url \@url }%
\providecommand \@url [1]{\endgroup\@href {#1}{\urlprefix }}%
\providecommand \urlprefix  [0]{URL }%
\providecommand \Eprint [0]{\href }%
\providecommand \doibase [0]{http://dx.doi.org/}%
\providecommand \selectlanguage [0]{\@gobble}%
\providecommand \bibinfo  [0]{\@secondoftwo}%
\providecommand \bibfield  [0]{\@secondoftwo}%
\providecommand \translation [1]{[#1]}%
\providecommand \BibitemOpen [0]{}%
\providecommand \bibitemStop [0]{}%
\providecommand \bibitemNoStop [0]{.\EOS\space}%
\providecommand \EOS [0]{\spacefactor3000\relax}%
\providecommand \BibitemShut  [1]{\csname bibitem#1\endcsname}%
\let\auto@bib@innerbib\@empty
%</preamble>
\bibitem [{\citenamefont {Carollo}\ \emph {et~al.}(2019)\citenamefont
  {Carollo}, \citenamefont {Spagnolo}, \citenamefont {Dubkov},\ and\
  \citenamefont {Valenti}}]{carollo19}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Angelo}\ \bibnamefont
  {Carollo}}, \bibinfo {author} {\bibfnamefont {Bernardo}\ \bibnamefont
  {Spagnolo}}, \bibinfo {author} {\bibfnamefont {Alexander~A.}\ \bibnamefont
  {Dubkov}}, \ and\ \bibinfo {author} {\bibfnamefont {Davide}\ \bibnamefont
  {Valenti}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {On quantumness
  in multi-parameter quantum estimation},}\ }\href {\doibase
  10.1088/1742-5468/ab3ccb} {\bibfield  {journal} {\bibinfo  {journal} {Journal
  of Statistical Mechanics: Theory and Experiment}\ }\textbf {\bibinfo {volume}
  {2019}},\ \bibinfo {pages} {094010} (\bibinfo {year} {2019})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Carollo}\ \emph {et~al.}(2020)\citenamefont
  {Carollo}, \citenamefont {Spagnolo}, \citenamefont {Dubkov},\ and\
  \citenamefont {Valenti}}]{carollo20}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Angelo}\ \bibnamefont
  {Carollo}}, \bibinfo {author} {\bibfnamefont {Bernardo}\ \bibnamefont
  {Spagnolo}}, \bibinfo {author} {\bibfnamefont {Alexander~A.}\ \bibnamefont
  {Dubkov}}, \ and\ \bibinfo {author} {\bibfnamefont {Davide}\ \bibnamefont
  {Valenti}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Erratum: {On}
  quantumness in multi-parameter quantum estimation (2019 {J}. {Stat}. {Mech}.
  094010)},}\ }\href {\doibase 10.1088/1742-5468/ab6f5e} {\bibfield  {journal}
  {\bibinfo  {journal} {Journal of Statistical Mechanics: Theory and
  Experiment}\ }\textbf {\bibinfo {volume} {2020}},\ \bibinfo {pages} {029902}
  (\bibinfo {year} {2020})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Tsang}\ \emph {et~al.}(2020)\citenamefont {Tsang},
  \citenamefont {Albarelli},\ and\ \citenamefont {Datta}}]{tsang20}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Mankei}\ \bibnamefont
  {Tsang}}, \bibinfo {author} {\bibfnamefont {Francesco}\ \bibnamefont
  {Albarelli}}, \ and\ \bibinfo {author} {\bibfnamefont {Animesh}\ \bibnamefont
  {Datta}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Quantum
  {Semiparametric} {Estimation}},}\ }\href {\doibase
  10.1103/PhysRevX.10.031023} {\bibfield  {journal} {\bibinfo  {journal}
  {Physical Review X}\ }\textbf {\bibinfo {volume} {10}},\ \bibinfo {pages}
  {031023} (\bibinfo {year} {2020})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Helstrom}(1976)}]{helstrom}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Carl~W.}\ \bibnamefont
  {Helstrom}},\ }\href
  {http://www.sciencedirect.com/science/bookseries/00765392/123} {\emph
  {\bibinfo {title} {Quantum Detection and Estimation Theory}}}\ (\bibinfo
  {publisher} {Academic Press},\ \bibinfo {address} {New York},\ \bibinfo
  {year} {1976})\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Hayashi}(2005)}]{hayashi05}%
  \BibitemOpen
  \bibinfo {editor} {\bibfnamefont {Masahito}\ \bibnamefont {Hayashi}},\ ed.,\
  \href {\doibase 10.1142/9789812563071} {\emph {\bibinfo {title} {Asymptotic
  Theory of Quantum Statistical Inference: Selected Papers}}}\ (\bibinfo
  {publisher} {World Scientific},\ \bibinfo {address} {Singapore},\ \bibinfo
  {year} {2005})\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Giovannetti}\ \emph {et~al.}(2011)\citenamefont
  {Giovannetti}, \citenamefont {Lloyd},\ and\ \citenamefont
  {Maccone}}]{glm2011}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Vittorio}\
  \bibnamefont {Giovannetti}}, \bibinfo {author} {\bibfnamefont {Seth}\
  \bibnamefont {Lloyd}}, \ and\ \bibinfo {author} {\bibfnamefont {Lorenzo}\
  \bibnamefont {Maccone}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Advances in quantum metrology},}\ }\href
  {http://dx.doi.org/10.1038/nphoton.2011.35} {\bibfield  {journal} {\bibinfo
  {journal} {Nature Photonics}\ }\textbf {\bibinfo {volume} {5}},\ \bibinfo
  {pages} {222--229} (\bibinfo {year} {2011})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Demkowicz-Dobrza{\'n}ski}\ \emph
  {et~al.}(2015)\citenamefont {Demkowicz-Dobrza{\'n}ski}, \citenamefont
  {Jarzyna},\ and\ \citenamefont {Ko{\l}ody{\'n}ski}}]{demkowicz15}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Rafa{\l}}\
  \bibnamefont {Demkowicz-Dobrza{\'n}ski}}, \bibinfo {author} {\bibfnamefont
  {Marcin}\ \bibnamefont {Jarzyna}}, \ and\ \bibinfo {author} {\bibfnamefont
  {Jan}\ \bibnamefont {Ko{\l}ody{\'n}ski}},\ }\bibfield  {title} {\enquote
  {\bibinfo {title} {{Quantum} {Limits} in {Optical} {Interferometry}},}\ }in\
  \href {\doibase 10.1016/bs.po.2015.02.003} {\emph {\bibinfo {booktitle}
  {Progress in {Optics}}}},\ Vol.~\bibinfo {volume} {60},\ \bibinfo {editor}
  {edited by\ \bibinfo {editor} {\bibfnamefont {E.}~\bibnamefont {Wolf}}}\
  (\bibinfo  {publisher} {Elsevier},\ \bibinfo {address} {Amsterdam},\ \bibinfo
  {year} {2015})\ Chap.~\bibinfo {chapter} {4}, pp.\ \bibinfo {pages}
  {345--435}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Pirandola}\ \emph {et~al.}(2018)\citenamefont
  {Pirandola}, \citenamefont {Bardhan}, \citenamefont {Gehring}, \citenamefont
  {Weedbrook},\ and\ \citenamefont {Lloyd}}]{pirandola18}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {S.}~\bibnamefont
  {Pirandola}}, \bibinfo {author} {\bibfnamefont {B.~R.}\ \bibnamefont
  {Bardhan}}, \bibinfo {author} {\bibfnamefont {T.}~\bibnamefont {Gehring}},
  \bibinfo {author} {\bibfnamefont {C.}~\bibnamefont {Weedbrook}}, \ and\
  \bibinfo {author} {\bibfnamefont {S.}~\bibnamefont {Lloyd}},\ }\bibfield
  {title} {\enquote {\bibinfo {title} {Advances in photonic quantum sensing},}\
  }\href {\doibase 10.1038/s41566-018-0301-6} {\bibfield  {journal} {\bibinfo
  {journal} {Nature Photonics}\ }\textbf {\bibinfo {volume} {12}},\ \bibinfo
  {pages} {724} (\bibinfo {year} {2018})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Pezz{\'e}}\ \emph {et~al.}(2018)\citenamefont
  {Pezz{\'e}}, \citenamefont {Smerzi}, \citenamefont {Oberthaler},
  \citenamefont {Schmied},\ and\ \citenamefont {Treutlein}}]{pezze18}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Luca}\ \bibnamefont
  {Pezz{\'e}}}, \bibinfo {author} {\bibfnamefont {Augusto}\ \bibnamefont
  {Smerzi}}, \bibinfo {author} {\bibfnamefont {Markus~K.}\ \bibnamefont
  {Oberthaler}}, \bibinfo {author} {\bibfnamefont {Roman}\ \bibnamefont
  {Schmied}}, \ and\ \bibinfo {author} {\bibfnamefont {Philipp}\ \bibnamefont
  {Treutlein}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Quantum
  metrology with nonclassical states of atomic ensembles},}\ }\href {\doibase
  10.1103/RevModPhys.90.035005} {\bibfield  {journal} {\bibinfo  {journal}
  {Reviews of Modern Physics}\ }\textbf {\bibinfo {volume} {90}},\ \bibinfo
  {pages} {035005} (\bibinfo {year} {2018})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Tsang}(2019{\natexlab{a}})}]{tsang19a}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Mankei}\ \bibnamefont
  {Tsang}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Resolving
  starlight: a quantum perspective},}\ }\href {\doibase
  10.1080/00107514.2020.1736375} {\bibfield  {journal} {\bibinfo  {journal}
  {Contemporary Physics}\ }\textbf {\bibinfo {volume} {60}},\ \bibinfo {pages}
  {279--298} (\bibinfo {year} {2019}{\natexlab{a}})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Holevo}(2011)}]{holevo11}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Alexander~S.}\
  \bibnamefont {Holevo}},\ }\href@noop {} {\emph {\bibinfo {title}
  {Probabilistic and Statistical Aspects of Quantum Theory}}}\ (\bibinfo
  {publisher} {Edizioni della Normale},\ \bibinfo {address} {Pisa, Italy},\
  \bibinfo {year} {2011})\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Personick}(1971)}]{personick71}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {S.}~\bibnamefont
  {Personick}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Application
  of quantum estimation theory to analog communication over quantum
  channels},}\ }\href {\doibase 10.1109/TIT.1971.1054643} {\bibfield  {journal}
  {\bibinfo  {journal} {IEEE Transactions on Information Theory}\ }\textbf
  {\bibinfo {volume} {17}},\ \bibinfo {pages} {240--246} (\bibinfo {year}
  {1971})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Yuen}\ and\ \citenamefont {Lax}(1973)}]{yuen_lax}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Horace~P.}\
  \bibnamefont {Yuen}}\ and\ \bibinfo {author} {\bibfnamefont {Melvin}\
  \bibnamefont {Lax}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Multiple-parameter quantum estimation and measurement of nonselfadjoint
  observables},}\ }\href {\doibase 10.1109/TIT.1973.1055103} {\bibfield
  {journal} {\bibinfo  {journal} {IEEE Transactions on Information Theory}\
  }\textbf {\bibinfo {volume} {19}},\ \bibinfo {pages} {740--750} (\bibinfo
  {year} {1973})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Tsuda}\ and\ \citenamefont
  {Matsumoto}(2005)}]{tsuda05}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Yoshiyuki}\
  \bibnamefont {Tsuda}}\ and\ \bibinfo {author} {\bibfnamefont {Keiji}\
  \bibnamefont {Matsumoto}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Quantum estimation for non-differentiable models},}\ }\href {\doibase
  10.1088/0305-4470/38/7/014} {\bibfield  {journal} {\bibinfo  {journal}
  {Journal of Physics A: Mathematical and General}\ }\textbf {\bibinfo {volume}
  {38}},\ \bibinfo {pages} {1593} (\bibinfo {year} {2005})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Tsang}(2012)}]{qzzb}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Mankei}\ \bibnamefont
  {Tsang}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Ziv-{Zakai}
  {Error} {Bounds} for {Quantum} {Parameter} {Estimation}},}\ }\href {\doibase
  10.1103/PhysRevLett.108.230401} {\bibfield  {journal} {\bibinfo  {journal}
  {Physical Review Letters}\ }\textbf {\bibinfo {volume} {108}},\ \bibinfo
  {pages} {230401} (\bibinfo {year} {2012})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Giovannetti}\ \emph {et~al.}(2012)\citenamefont
  {Giovannetti}, \citenamefont {Lloyd},\ and\ \citenamefont
  {Maccone}}]{glm2012}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Vittorio}\
  \bibnamefont {Giovannetti}}, \bibinfo {author} {\bibfnamefont {Seth}\
  \bibnamefont {Lloyd}}, \ and\ \bibinfo {author} {\bibfnamefont {Lorenzo}\
  \bibnamefont {Maccone}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Quantum measurement bounds beyond the uncertainty relations},}\ }\href
  {\doibase 10.1103/PhysRevLett.108.260405} {\bibfield  {journal} {\bibinfo
  {journal} {Physical Review Letters}\ }\textbf {\bibinfo {volume} {108}},\
  \bibinfo {pages} {260405} (\bibinfo {year} {2012})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Hall}\ and\ \citenamefont
  {Wiseman}(2012)}]{hall_prx}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Michael J.~W.}\
  \bibnamefont {Hall}}\ and\ \bibinfo {author} {\bibfnamefont {Howard~M.}\
  \bibnamefont {Wiseman}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {{Does Nonlinear Metrology Offer Improved Resolution? Answers from Quantum
  Information Theory}},}\ }\href {\doibase 10.1103/PhysRevX.2.041006}
  {\bibfield  {journal} {\bibinfo  {journal} {Physical Review X}\ }\textbf
  {\bibinfo {volume} {2}},\ \bibinfo {pages} {041006} (\bibinfo {year}
  {2012})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Berry}\ \emph {et~al.}(2015)\citenamefont {Berry},
  \citenamefont {Tsang}, \citenamefont {Hall},\ and\ \citenamefont
  {Wiseman}}]{qbzzb}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Dominic~W.}\
  \bibnamefont {Berry}}, \bibinfo {author} {\bibfnamefont {Mankei}\
  \bibnamefont {Tsang}}, \bibinfo {author} {\bibfnamefont {Michael J.~W.}\
  \bibnamefont {Hall}}, \ and\ \bibinfo {author} {\bibfnamefont {Howard~M.}\
  \bibnamefont {Wiseman}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Quantum {Bell}-{Ziv}-{Zakai} {Bounds} and {Heisenberg} {Limits} for
  {Waveform} {Estimation}},}\ }\href {\doibase 10.1103/PhysRevX.5.031018}
  {\bibfield  {journal} {\bibinfo  {journal} {Physical Review X}\ }\textbf
  {\bibinfo {volume} {5}},\ \bibinfo {pages} {031018} (\bibinfo {year}
  {2015})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Lu}\ and\ \citenamefont {Tsang}(2016)}]{qwwb}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Xiao-Ming}\
  \bibnamefont {Lu}}\ and\ \bibinfo {author} {\bibfnamefont {Mankei}\
  \bibnamefont {Tsang}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Quantum {Weiss}-{Weinstein} bounds for quantum metrology},}\ }\href
  {\doibase 10.1088/2058-9565/1/1/015002} {\bibfield  {journal} {\bibinfo
  {journal} {Quantum Science and Technology}\ }\textbf {\bibinfo {volume}
  {1}},\ \bibinfo {pages} {015002} (\bibinfo {year} {2016})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Nair}(2018)}]{nair18}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Ranjith}\ \bibnamefont
  {Nair}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Fundamental
  quantum limits in optical metrology from rate-distortion theory},}\ }\href
  {\doibase 10.1088/1751-8121/aade9f} {\bibfield  {journal} {\bibinfo
  {journal} {Journal of Physics A: Mathematical and Theoretical}\ }\textbf
  {\bibinfo {volume} {51}},\ \bibinfo {pages} {434001} (\bibinfo {year}
  {2018})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Rubio}\ and\ \citenamefont
  {Dunningham}(2019)}]{rubio19}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Jes{\'u}s}\
  \bibnamefont {Rubio}}\ and\ \bibinfo {author} {\bibfnamefont {Jacob}\
  \bibnamefont {Dunningham}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Quantum metrology in the presence of limited data},}\ }\href {\doibase
  10.1088/1367-2630/ab098b} {\bibfield  {journal} {\bibinfo  {journal} {New
  Journal of Physics}\ }\textbf {\bibinfo {volume} {21}},\ \bibinfo {pages}
  {043037} (\bibinfo {year} {2019})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Rubio}\ and\ \citenamefont
  {Dunningham}(2020)}]{rubio20}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Jesús}\ \bibnamefont
  {Rubio}}\ and\ \bibinfo {author} {\bibfnamefont {Jacob}\ \bibnamefont
  {Dunningham}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Bayesian
  multiparameter quantum metrology with limited data},}\ }\href {\doibase
  10.1103/PhysRevA.101.032114} {\bibfield  {journal} {\bibinfo  {journal}
  {Physical Review A}\ }\textbf {\bibinfo {volume} {101}},\ \bibinfo {pages}
  {032114} (\bibinfo {year} {2020})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Gill}\ and\ \citenamefont {{Gu{\c t}{\u
  a}}}(2013)}]{gill_guta}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Richard~D.}\
  \bibnamefont {Gill}}\ and\ \bibinfo {author} {\bibfnamefont {M{\u a}d{\u
  a}lin}\ \bibnamefont {{Gu{\c t}{\u a}}}},\ }\enquote {\bibinfo {title} {On
  asymptotic quantum statistical inference},}\ in\ \href {\doibase
  10.1214/12-IMSCOLL909} {\emph {\bibinfo {booktitle} {From Probability to
  Statistics and Back: High-Dimensional Models and Processes -- A Festschrift
  in Honor of Jon A. Wellner}}},\ \bibinfo {series} {Collections},
  Vol.~\bibinfo {volume} {9},\ \bibinfo {editor} {edited by\ \bibinfo {editor}
  {\bibfnamefont {M.}~\bibnamefont {Banerjee}}, \bibinfo {editor}
  {\bibfnamefont {F.}~\bibnamefont {Bunea}}, \bibinfo {editor} {\bibfnamefont
  {J.}~\bibnamefont {Huang}}, \bibinfo {editor} {\bibfnamefont
  {V.}~\bibnamefont {Koltchinskii}}, \ and\ \bibinfo {editor} {\bibfnamefont
  {M.~H.}\ \bibnamefont {Maathuis}}}\ (\bibinfo  {publisher} {Institute of
  Mathematical Statistics},\ \bibinfo {address} {Beachwood, Ohio, USA},\
  \bibinfo {year} {2013})\ pp.\ \bibinfo {pages} {105--127}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Yamagata}\ \emph {et~al.}(2013)\citenamefont
  {Yamagata}, \citenamefont {Fujiwara},\ and\ \citenamefont
  {Gill}}]{yamagata13}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Koichi}\ \bibnamefont
  {Yamagata}}, \bibinfo {author} {\bibfnamefont {Akio}\ \bibnamefont
  {Fujiwara}}, \ and\ \bibinfo {author} {\bibfnamefont {Richard~D.}\
  \bibnamefont {Gill}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Quantum local asymptotic normality based on a new quantum likelihood
  ratio},}\ }\href {\doibase 10.1214/13-AOS1147} {\bibfield  {journal}
  {\bibinfo  {journal} {The Annals of Statistics}\ }\textbf {\bibinfo {volume}
  {41}},\ \bibinfo {pages} {2197--2217} (\bibinfo {year} {2013})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Yang}\ \emph {et~al.}(2019)\citenamefont {Yang},
  \citenamefont {Chiribella},\ and\ \citenamefont {Hayashi}}]{yang19}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Yuxiang}\ \bibnamefont
  {Yang}}, \bibinfo {author} {\bibfnamefont {Giulio}\ \bibnamefont
  {Chiribella}}, \ and\ \bibinfo {author} {\bibfnamefont {Masahito}\
  \bibnamefont {Hayashi}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Attaining the {Ultimate} {Precision} {Limit} in {Quantum} {State}
  {Estimation}},}\ }\href {\doibase 10.1007/s00220-019-03433-4} {\bibfield
  {journal} {\bibinfo  {journal} {Communications in Mathematical Physics}\
  }\textbf {\bibinfo {volume} {368}},\ \bibinfo {pages} {223--293} (\bibinfo
  {year} {2019})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Matsumoto}(2002)}]{matsumoto02}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {K.}~\bibnamefont
  {Matsumoto}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {A new
  approach to the {Cram\'er}-{Rao}-type bound of the pure-state model},}\
  }\href {\doibase 10.1088/0305-4470/35/13/307} {\bibfield  {journal} {\bibinfo
   {journal} {Journal of Physics A: Mathematical and General}\ }\textbf
  {\bibinfo {volume} {35}},\ \bibinfo {pages} {3111--3123} (\bibinfo {year}
  {2002})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Ragy}\ \emph {et~al.}(2016)\citenamefont {Ragy},
  \citenamefont {Jarzyna},\ and\ \citenamefont
  {Demkowicz-Dobrzański}}]{ragy16}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Sammy}\ \bibnamefont
  {Ragy}}, \bibinfo {author} {\bibfnamefont {Marcin}\ \bibnamefont {Jarzyna}},
  \ and\ \bibinfo {author} {\bibfnamefont {Rafał}\ \bibnamefont
  {Demkowicz-Dobrzański}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Compatibility in multiparameter quantum metrology},}\ }\href {\doibase
  10.1103/PhysRevA.94.052108} {\bibfield  {journal} {\bibinfo  {journal}
  {Physical Review A}\ }\textbf {\bibinfo {volume} {94}},\ \bibinfo {pages}
  {052108} (\bibinfo {year} {2016})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Szczykulska}\ \emph {et~al.}(2016)\citenamefont
  {Szczykulska}, \citenamefont {Baumgratz},\ and\ \citenamefont
  {Datta}}]{szczykulska16}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Magdalena}\
  \bibnamefont {Szczykulska}}, \bibinfo {author} {\bibfnamefont {Tillmann}\
  \bibnamefont {Baumgratz}}, \ and\ \bibinfo {author} {\bibfnamefont {Animesh}\
  \bibnamefont {Datta}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Multi-parameter quantum metrology},}\ }\href {\doibase
  10.1080/23746149.2016.1230476} {\bibfield  {journal} {\bibinfo  {journal}
  {Advances in Physics: X}\ }\textbf {\bibinfo {volume} {1}},\ \bibinfo {pages}
  {621--639} (\bibinfo {year} {2016})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Bradshaw}\ \emph {et~al.}(2017)\citenamefont
  {Bradshaw}, \citenamefont {Assad},\ and\ \citenamefont {Lam}}]{bradshaw17}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Mark}\ \bibnamefont
  {Bradshaw}}, \bibinfo {author} {\bibfnamefont {Syed~M.}\ \bibnamefont
  {Assad}}, \ and\ \bibinfo {author} {\bibfnamefont {Ping~Koy}\ \bibnamefont
  {Lam}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {A tight
  {Cram\'er}–{Rao} bound for joint parameter estimation with a pure two-mode
  squeezed probe},}\ }\href {\doibase 10.1016/j.physleta.2017.06.024}
  {\bibfield  {journal} {\bibinfo  {journal} {Physics Letters A}\ }\textbf
  {\bibinfo {volume} {381}},\ \bibinfo {pages} {2598--2607} (\bibinfo {year}
  {2017})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Bradshaw}\ \emph {et~al.}(2018)\citenamefont
  {Bradshaw}, \citenamefont {Lam},\ and\ \citenamefont {Assad}}]{bradshaw18}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Mark}\ \bibnamefont
  {Bradshaw}}, \bibinfo {author} {\bibfnamefont {Ping~Koy}\ \bibnamefont
  {Lam}}, \ and\ \bibinfo {author} {\bibfnamefont {Syed~M.}\ \bibnamefont
  {Assad}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Ultimate
  precision of joint quadrature parameter estimation with a {Gaussian}
  probe},}\ }\href {\doibase 10.1103/PhysRevA.97.012106} {\bibfield  {journal}
  {\bibinfo  {journal} {Physical Review A}\ }\textbf {\bibinfo {volume} {97}},\
  \bibinfo {pages} {012106} (\bibinfo {year} {2018})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Albarelli}\ \emph {et~al.}(2019)\citenamefont
  {Albarelli}, \citenamefont {Friel},\ and\ \citenamefont
  {Datta}}]{albarelli19}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Francesco}\
  \bibnamefont {Albarelli}}, \bibinfo {author} {\bibfnamefont {Jamie~F.}\
  \bibnamefont {Friel}}, \ and\ \bibinfo {author} {\bibfnamefont {Animesh}\
  \bibnamefont {Datta}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Evaluating the {Holevo} {Cram\'er}-{Rao} {Bound} for {Multiparameter}
  {Quantum} {Metrology}},}\ }\href {\doibase 10.1103/PhysRevLett.123.200503}
  {\bibfield  {journal} {\bibinfo  {journal} {Physical Review Letters}\
  }\textbf {\bibinfo {volume} {123}},\ \bibinfo {pages} {200503} (\bibinfo
  {year} {2019})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Suzuki}\ \emph {et~al.}(2020)\citenamefont {Suzuki},
  \citenamefont {Yang},\ and\ \citenamefont {Hayashi}}]{suzuki19}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Jun}\ \bibnamefont
  {Suzuki}}, \bibinfo {author} {\bibfnamefont {Yuxiang}\ \bibnamefont {Yang}},
  \ and\ \bibinfo {author} {\bibfnamefont {Masahito}\ \bibnamefont {Hayashi}},\
  }\bibfield  {title} {\enquote {\bibinfo {title} {Quantum state estimation
  with nuisance parameters},}\ }\href {http://arxiv.org/abs/1911.02790}
  {\bibfield  {journal} {\bibinfo  {journal} {arXiv:1911.02790 [quant-ph]}\ }
  (\bibinfo {year} {2020})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Hayashi}(2017)}]{hayashi}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Masahito}\
  \bibnamefont {Hayashi}},\ }\href {\doibase 10.1007/978-3-662-49725-8} {\emph
  {\bibinfo {title} {Quantum {I}nformation {T}heory: {M}athematical
  {F}oundation}}},\ \bibinfo {edition} {2nd}\ ed.\ (\bibinfo  {publisher}
  {Springer},\ \bibinfo {address} {Berlin},\ \bibinfo {year}
  {2017})\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Paris}(2009)}]{paris}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Matteo G.~A.}\
  \bibnamefont {Paris}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Quantum estimation for quantum technology},}\ }\href {\doibase
  10.1142/S0219749909004839} {\bibfield  {journal} {\bibinfo  {journal}
  {International Journal of Quantum Information}\ }\textbf {\bibinfo {volume}
  {07}},\ \bibinfo {pages} {125--137} (\bibinfo {year} {2009})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Escher}\ \emph {et~al.}(2011)\citenamefont {Escher},
  \citenamefont {de~Matos~Filho},\ and\ \citenamefont {Davidovich}}]{escher}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {B.~M.}\ \bibnamefont
  {Escher}}, \bibinfo {author} {\bibfnamefont {R.~L.}\ \bibnamefont
  {de~Matos~Filho}}, \ and\ \bibinfo {author} {\bibfnamefont {L.}~\bibnamefont
  {Davidovich}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {General
  framework for estimating the ultimate precision limit in noisy
  quantum-enhanced metrology},}\ }\href {\doibase 10.1038/nphys1958} {\bibfield
   {journal} {\bibinfo  {journal} {Nature Physics}\ }\textbf {\bibinfo {volume}
  {7}},\ \bibinfo {pages} {406--411} (\bibinfo {year} {2011})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Tsang}\ \emph {et~al.}(2011)\citenamefont {Tsang},
  \citenamefont {Wiseman},\ and\ \citenamefont {Caves}}]{twc}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Mankei}\ \bibnamefont
  {Tsang}}, \bibinfo {author} {\bibfnamefont {Howard~M.}\ \bibnamefont
  {Wiseman}}, \ and\ \bibinfo {author} {\bibfnamefont {Carlton~M.}\
  \bibnamefont {Caves}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Fundamental quantum limit to waveform estimation},}\ }\href {\doibase
  10.1103/PhysRevLett.106.090401} {\bibfield  {journal} {\bibinfo  {journal}
  {Physical Review Letters}\ }\textbf {\bibinfo {volume} {106}},\ \bibinfo
  {pages} {090401} (\bibinfo {year} {2011})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {{Gu{\c t}{\u a}}}(2011)}]{guta11}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {M{\u a}d{\u a}lin}\
  \bibnamefont {{Gu{\c t}{\u a}}}},\ }\bibfield  {title} {\enquote {\bibinfo
  {title} {Fisher information and asymptotic normality in system identification
  for quantum {Markov} chains},}\ }\href {\doibase 10.1103/PhysRevA.83.062324}
  {\bibfield  {journal} {\bibinfo  {journal} {Physical Review A}\ }\textbf
  {\bibinfo {volume} {83}},\ \bibinfo {pages} {062324} (\bibinfo {year}
  {2011})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Tsang}(2013)}]{tsang_open}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Mankei}\ \bibnamefont
  {Tsang}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Quantum metrology
  with open dynamical systems},}\ }\href {\doibase
  10.1088/1367-2630/15/7/073005} {\bibfield  {journal} {\bibinfo  {journal}
  {New Journal of Physics}\ }\textbf {\bibinfo {volume} {15}},\ \bibinfo
  {pages} {073005} (\bibinfo {year} {2013})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Alipour}\ and\ \citenamefont
  {Rezakhani}(2015)}]{alipour}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {S.}~\bibnamefont
  {Alipour}}\ and\ \bibinfo {author} {\bibfnamefont {A.~T.}\ \bibnamefont
  {Rezakhani}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Extended
  convexity of quantum fisher information in quantum metrology},}\ }\href
  {\doibase 10.1103/PhysRevA.91.042104} {\bibfield  {journal} {\bibinfo
  {journal} {Physical Review A}\ }\textbf {\bibinfo {volume} {91}},\ \bibinfo
  {pages} {042104} (\bibinfo {year} {2015})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Ng}\ \emph {et~al.}(2016)\citenamefont {Ng},
  \citenamefont {Ang}, \citenamefont {Wheatley}, \citenamefont {Yonezawa},
  \citenamefont {Furusawa}, \citenamefont {Huntington},\ and\ \citenamefont
  {Tsang}}]{ng16}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Shilin}\ \bibnamefont
  {Ng}}, \bibinfo {author} {\bibfnamefont {Shan~Zheng}\ \bibnamefont {Ang}},
  \bibinfo {author} {\bibfnamefont {Trevor~A.}\ \bibnamefont {Wheatley}},
  \bibinfo {author} {\bibfnamefont {Hidehiro}\ \bibnamefont {Yonezawa}},
  \bibinfo {author} {\bibfnamefont {Akira}\ \bibnamefont {Furusawa}}, \bibinfo
  {author} {\bibfnamefont {Elanor~H.}\ \bibnamefont {Huntington}}, \ and\
  \bibinfo {author} {\bibfnamefont {Mankei}\ \bibnamefont {Tsang}},\ }\bibfield
   {title} {\enquote {\bibinfo {title} {Spectrum analysis with quantum
  dynamical systems},}\ }\href {\doibase 10.1103/PhysRevA.93.042121} {\bibfield
   {journal} {\bibinfo  {journal} {Physical Review A}\ }\textbf {\bibinfo
  {volume} {93}},\ \bibinfo {pages} {042121} (\bibinfo {year}
  {2016})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Yuan}\ and\ \citenamefont {Fung}(2017)}]{yuan17a}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Haidong}\ \bibnamefont
  {Yuan}}\ and\ \bibinfo {author} {\bibfnamefont {Chi-Hang~Fred}\ \bibnamefont
  {Fung}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Quantum metrology
  matrix},}\ }\href {\doibase 10.1103/PhysRevA.96.012310} {\bibfield  {journal}
  {\bibinfo  {journal} {Physical Review A}\ }\textbf {\bibinfo {volume} {96}},\
  \bibinfo {pages} {012310} (\bibinfo {year} {2017})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Sidhu}\ and\ \citenamefont {Kok}(2020)}]{sidhu20}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Jasminder~S.}\
  \bibnamefont {Sidhu}}\ and\ \bibinfo {author} {\bibfnamefont {Pieter}\
  \bibnamefont {Kok}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {{Geometric} {perspective} on {quantum} {parameter} {estimation}},}\ }\href
  {\doibase 10.1116/1.5119961} {\bibfield  {journal} {\bibinfo  {journal} {AVS
  Quantum Science}\ }\textbf {\bibinfo {volume} {2}},\ \bibinfo {pages}
  {014701} (\bibinfo {year} {2020})}\BibitemShut {NoStop}%
\bibitem [{\citenamefont {Genoni}\ and\ \citenamefont
  {Tufarelli}(2019)}]{genoni19}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Marco~G.}\
  \bibnamefont {Genoni}}\ and\ \bibinfo {author} {\bibfnamefont {Tommaso}\
  \bibnamefont {Tufarelli}},\ }\bibfield  {title} {\enquote {\bibinfo {title}
  {Non-orthogonal bases for quantum metrology},}\ }\href {\doibase
  10.1088/1751-8121/ab3fe0} {\bibfield  {journal} {\bibinfo  {journal} {Journal
  of Physics A: Mathematical and Theoretical}\ }\textbf {\bibinfo {volume}
  {52}},\ \bibinfo {pages} {434002} (\bibinfo {year} {2019})}\BibitemShut
  {NoStop}%
\bibitem [{\citenamefont {Tsang}(2019{\natexlab{b}})}]{tsang19c}%
  \BibitemOpen
  \bibfield  {author} {\bibinfo {author} {\bibfnamefont {Mankei}\ \bibnamefont
  {Tsang}},\ }\bibfield  {title} {\enquote {\bibinfo {title} {Quantum analogs
  of the conditional expectation for retrodiction and smoothing: a unified
  view},}\ }\href {http://arxiv.org/abs/1912.02711} {\bibfield  {journal}
  {\bibinfo  {journal} {arXiv:1912.02711 [quant-ph]}\ } (\bibinfo {year}
  {2019}{\natexlab{b}})}\BibitemShut {NoStop}%
\end{thebibliography}%

\documentclass[aps,prl,10pt,amsmath,twocolumn,floatfix,longbibliography]{revtex4-1}
\usepackage[dvipsnames]{xcolor}
\usepackage[colorlinks=true,bookmarks=false,linkcolor=NavyBlue,urlcolor=NavyBlue,citecolor=NavyBlue,breaklinks]{hyperref}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{tabularx}
\usepackage{graphicx}
\usepackage{times}
\usepackage{mathtools}
\usepackage{braket}
\usepackage{amsthm}
 
\newtheorem{theorem}{Theorem}

\newtheorem{proposition}{Proposition}

\theoremstyle{remark}

\newcommand{\parti}[2]{\frac{\partial #1}{\partial #2}}

\newcommand{\abs}[1]{\left|#1\right|}
\newcommand{\bk}[1]{\left(#1\right)}
\newcommand{\Bk}[1]{\left[#1\right]}
\newcommand{\BK}[1]{\left\{#1\right\}}

\newcommand{\norm}[1]{\lVert #1 \rVert}

\DeclareMathOperator{\trace}{tr}
\DeclareMathOperator{\real}{Re}
\DeclareMathOperator{\imag}{Im}

\begin{document}

\title{The Holevo Cram\'er-Rao bound is at most
thrice the Helstrom version}

\author{Mankei Tsang}
\email{mankei@nus.edu.sg}
\homepage{https://blog.nus.edu.sg/mankei/}
\affiliation{Department of Electrical and Computer Engineering,
  National University of Singapore, 4 Engineering Drive 3, Singapore
  117583}

\affiliation{Department of Physics, National University of Singapore,
  2 Science Drive 3, Singapore 117551}

\date{\today}


\begin{abstract}
  In quantum metrology, the Holevo Cram\'er-Rao bound has attracted
  renewed interest in recent years due to its superiority over the
  Helstrom Cram\'er-Rao bound and its asymptotic attainability for
  multiparameter estimation. Its evaluation, however, is often much
  more difficult than that of the Helstrom version, calling into
  question the actual improvement offered by the Holevo bound and
  whether it is worth the trouble. Here I prove that the Holevo bound
  is at most thrice the Helstrom version, so the improvement must be
  limited and the role of incompatibility in quantum estimation turns
  out to be modest. The result also shows that the Helstrom version
  remains a pretty good bound even for multiple parameters, as it can
  be approached asymptotically to within a factor of 3.

  \textbf{Update}: References~\cite{carollo19,carollo20,tsang20}
  supersede this work by proving that the Holevo Cram\'er-Rao bound
  is, in fact, at most twice the Helstrom version and the factor-of-2
  bound can be tight.
\end{abstract}

\maketitle

For any measurement of a quantum system and any unbiased estimator, a
quantum generalization of the Cram\'er-Rao bound (CRB)---first
proposed by Helstrom in 1967 \cite{helstrom}---can be expressed as
\cite{hayashi05}
\begin{align}
\trace G \Sigma &\ge C^{S} \equiv \min_{X \in \mathcal X} \trace \real Q(X),
\label{helstrom}
\\
Q(X) &\equiv \sqrt{G} Z(X)\sqrt{G},
\quad
Z_{\mu\nu}(X) \equiv \trace \rho X_\mu X_\nu,
\end{align}
where $\Sigma$ is the error covariance matrix, $G$ is a real
 and positive-semidefinite cost matrix, $\rho$ is the
density operator of the quantum system that depends on $n$ real
unknown parameters $\theta = (\theta_1,\theta_2,\dots,\theta_n)$,
$\mathcal X$ is the set of all vectoral Hermitian operators
$X = (X_1,X_2,\dots,X_n)$ that satisfy
$\trace X_\mu\partial \rho/\partial \theta_\nu = \delta_{\mu\nu}$, and
the real part of a matrix is defined by
$(\real Q)_{\mu\nu} = \real(Q_{\mu\nu}) =
[Q_{\mu\nu}+(Q_{\mu\nu})^*]/2$.  The original form of $C^{S}$ in terms
of the symmetric logarithmic derivatives of $\rho$
\cite{helstrom,hayashi05} is a closed-form solution of
Eq.~(\ref{helstrom}). The Helstrom CRB serves as a fundamental limit
to quantum estimation and has found many applications in quantum
metrology \cite{glm2011,demkowicz15,pirandola18,pezze18,tsang19a}.

Despite the popularity of the Helstrom CRB, better bounds exist
\cite{holevo11,personick71,yuen_lax,tsuda05,qzzb,glm2012,hall_prx,qbzzb,qwwb,nair18,rubio19,
  rubio20}. In particular, Holevo proposed a bound that can be
expressed as \cite{holevo11,hayashi05,gill_guta,yamagata13}
\begin{align}
  \trace G \Sigma &\ge C^{H} \ge \max\BK{C^{S},C^{R}},
  \\
  C^{H} &\equiv \min_{X \in \mathcal X} \Bk{\trace\real Q(X)+ \norm{\imag Q(X)}_1},
\label{hcrb}
\end{align}
where $C^{R}$ is another CRB due to Yuen and Lax \cite{yuen_lax} that
is not elaborated here, the imaginary part of a matrix is defined by
$(\imag Q)_{\mu\nu} = \imag(Q_{\mu\nu}) =
[Q_{\mu\nu}-(Q_{\mu\nu})^*]/(2i)$, the trace norm is defined as
$\norm{A}_1 \equiv \trace \sqrt{A^\dagger A}$, and $\dagger$ denotes
the conjugate transpose.  When there are multiple parameters, the
Holevo CRB $C^{H}$ is not only tighter but also attainable
asymptotically \cite{gill_guta,yamagata13,yang19}, as it accounts
properly for any incompatibility of the observables that should be
measured. The bound has attracted renewed interest in recent years
\cite{gill_guta,matsumoto02,hayashi05,yamagata13,ragy16,szczykulska16,bradshaw17,bradshaw18,yang19,albarelli19,suzuki19},
as many applications involve multiple unknown parameters and the
effect of incompatibility is of both fundamental and practical
interest.

Despite the fundamental importance of the Holevo CRB, its evaluation
is difficult and daunting numerics is often needed.  This is in
contrast to the more amenable Helstrom CRB, for which many fruitful
computation techniques have been devised over the years
\cite{helstrom,hayashi05,glm2011,demkowicz15,pirandola18,pezze18,tsang19a,hayashi,paris,escher,twc,guta11,tsang_open,alipour,ng16,yuan17a,sidhu20,genoni19,tsang19c}. For
researchers who are reluctant to undertake the endeavor, this raises
the questions how much improvement the Holevo CRB can actually offer
and whether it is worth the trouble after all.  The following theorem
gives a concrete answer.
\begin{theorem}
$C^{H} \le 3 C^{S}$.
\label{thrice}
\end{theorem}

\begin{proof}
  For any $X$, it can be shown that $\sqrt{G}$, $Z$, $Q$, and
  $\real Q$ are positive-semidefinite, $\sqrt{G}$, $\real Q$, and
  $\imag Q$ are real, $i\imag Q$ is Hermitian, and $\imag Q$ is
  skew-symmetric. With
\begin{align}
Q &= \real Q + i \imag Q,
&
i\imag Q &= Q - \real Q,
\end{align}
one can derive an uncertainty relation given by
\begin{align}
\norm{\imag Q}_1 &= \norm{i\imag Q}_1
= \norm{Q - \real Q}_1
\le \norm{Q}_1 + \norm{\real Q}_1
\nonumber\\
&= \trace Q + \trace \real Q = 2 \trace \real Q,
\label{ineq}
\end{align}
where the triangle inequality is used, $\norm{Q}_1 = \trace Q$ and
$\norm{\real Q}_1 = \trace \real Q$ because $Q$ and $\real Q$ are
positive-semidefinite, and
$\trace Q = \trace \real Q + i \trace \imag Q = \trace \real Q$
because $\imag Q$ is skew-symmetric. Now write the Helstrom CRB as
\begin{align}
C^{S} &= \trace \real Q(X^{S}), 
\label{helstrom2}
\end{align}
where $X^{S}$ is the element in $\mathcal X$ that minimizes
$\trace \real Q(X)$ in Eq.~(\ref{helstrom}).  Combining
Eqs.~(\ref{hcrb}), (\ref{ineq}), and (\ref{helstrom2}), one obtains
\begin{align}
C^{H} &\le \trace\real Q(X^{S}) + \norm{\imag Q(X^{S})}_1
\label{CH_bound}
\\
&\le 3 \trace\real Q(X^{S}) = 3 C^{S}.
\end{align}
\end{proof}

Theorem~\ref{thrice} puts the Holevo CRB in the sandwich
\begin{align}
\max\BK{C^{S},C^{R}} \le C^{H} \le 3C^{S},
\end{align}
and researchers can now decide for themselves whether an improvement
by at most a factor of 3 warrants the extra effort of evaluating
$C^{H}$. The theorem may even be on the generous side, as numerical
results often show that the improvement is less than a factor of 2
\cite{bradshaw17,bradshaw18,albarelli19}. On the flip side,
Theorem~\ref{thrice}, together with the asymptotic attainability of
$C^{H}$ \cite{gill_guta,yamagata13,yang19}, implies that $C^{S}$ is
asymptotically approachable to within a factor of 3, so the Helstrom
CRB turns out to be a pretty good bound after all, even for multiple
parameters.

The bound in Theorem~\ref{thrice} can be further tightened in special
cases. Here I consider the cases where $G$ is rank-one or rank-two.
\begin{proposition}
If $G$ is rank-one, $C^H = C^S$.
\label{rankone}
\end{proposition}

\begin{proof}
 A rank-one $G$ can be expressed as $G = g e e^\top$,
where $g$ is its real and positive eigenvalue,
$e$ is the real unit eigenvector, and $\top$ denotes
the transpose. Then 
\begin{align}
\sqrt{G} &= \sqrt{g} e e^\top,
&
Q &=\sqrt{G} Z \sqrt{G} =  g(e^\top Z e) e e^\top.
\end{align}
Since $Z \ge 0$, $e^\top Z e$ is real and nonnegative, meaning that
$Q$ is real and $\imag Q = 0$. The $C^H$ given by Eq.~(\ref{hcrb}) is
hence equal to the $C^S$ given by Eq.~(\ref{helstrom}).
\end{proof}

Note that a rank-one $G$ is not the same as the case of $n = 1$
unknown parameter. To be specific, let $\beta(\theta)$ be a scalar
parameter of interest that depends on the $n$ unknown parameters
$\theta$.  For example, if $\beta(\theta) = \theta_1$, then the rest
$(\theta_2,\dots,\theta_n)$ are nuisance parameters, which may hamper
the estimation. A bound on the error of estimating $\beta$ in the
presence of the many unknowns can be obtained by assuming
\begin{align}
G_{\mu\nu} &= \parti{\beta}{\theta_\mu}\parti{\beta}{\theta_\nu}.
\end{align}
It follows that
\begin{align}
Q &= \bk{\trace \rho Y^2} e e^\top,
\quad
Y = \sum_{\mu=1}^n \parti{\beta}{\theta_\mu}X_\mu,
\\
C^H &= C^S = \min_{Y \in \mathcal Y} \trace \rho Y^2,
\end{align}
where $\mathcal Y$ is the set of all Hermitian operators that satisfy
the constraints
\begin{align}
\trace Y \parti{\rho}{\theta_\mu} &= \parti{\beta}{\theta_\mu},
\quad \mu = 1,2,\dots,n.
\end{align}
This formulation of $C^S$ is then equivalent to the one in
Ref.~\cite{tsang19c} for semiparametric estimation.  For the problems
studied in Ref.~\cite{tsang19c}, Proposition~\ref{rankone} implies
that the Holevo CRB offers no improvement and the Helstrom CRBs
computed there are asymptotically attainable, at least when $n$ is
finite.

\begin{proposition}
If $G$ is rank-two, $C^H \le 2 C^S$.
\label{ranktwo}
\end{proposition}

\begin{proof}
  Let the positive eigenvalues of a rank-$r$ $G$ be
  $\{g_j:j = 1,2,\dots,r\}$ and the corresponding real unit
  eigenvectors be $\{e^j: j = 1,2,\dots,r\}$. Then
\begin{align}
Q &= \sum_{j=1}^r\sum_{k=1}^r \bk{\trace \rho Y_j Y_k}e^j e^{k\top},
&
Y_j &= \sqrt{g_j} \sum_{\mu=1}^n e^j_{\mu} X_\mu.
\end{align}
If $r = 2$,
\begin{align}
\imag Q &= (\imag \trace \rho Y_1Y_2) \bk{e^1 e^{2\top}- e^2 e^{1\top}},
\\
\norm{\imag Q}_1 &= 2\abs{\imag \trace \rho Y_1Y_2} 
= \abs{i\trace \rho \Bk{Y_1,Y_2}}
\\
&\le \trace \rho Y_1^2 + \trace \rho Y_2^2 = \trace \real Q,
\end{align}
where the inequality comes from
Ref.~\cite[Proposition~2.8.3]{holevo11}.  Hence
\begin{align}
C^{H} &\le \trace\real Q(X^{S}) + \norm{\imag Q(X^{S})}_1
\\
&\le 2 \trace\real Q(X^{S}) = 2 C^{S}.
\end{align}
\end{proof}
Considering only the propositions, one might suspect that $C^H$ could
become significantly higher for a $G$ with a higher rank, but
Theorem~\ref{thrice} settles the general case by imposing a hard limit
for any rank, revealing the surprisingly modest role of
incompatibility in asymptotic quantum estimation.  It remains an open
question whether Theorem~\ref{thrice} can be improved further and a
tighter upper bound on $C^H$ can be found.

Discussions with Francesco Albarelli, Richard Gill, and Madalin Guta
are gratefully acknowledged. This work is supported by the Singapore
National Research Foundation under Project No.~QEP-P7.
\bibliography{research2}

\end{document}